# Comparative study of magnetic exchange parameters and magnon dispersions in NiO and MnO from first principles

### v1

@@ INTRODUCTION
Magnetic materials play a pivotal role in a wide range of modern technologies, from data storage and electronic devices to renewable energy solutions. They hold further promises in spintronics, magnonics, and quantum information 1, 2. Noteworthy classes of magnetic materials that have drawn attention include high-temperature superconductors 3, 4, multiferroics 5, 6, 2D magnetic materials 7, 8, magnetic topological materials 9, 10, and, recently, altermagnets 11, 12, 13. Various experimental techniques, such as inelastic neutron scattering, resonant inelastic X-ray scattering, and spin-polarized electron energy loss spectroscopy are employed to study these materials. Yet, theoretical models are essential for the interpretation of these experiments. Spin model Hamiltonians, such as the Ising, XY, and Heisenberg models 14 are powerful tools for understanding the microscopic origin of macroscopic properties and describing collective spin excitations and their coupling to phonons 15 and plasmons 16. The Heisenberg model typically incorporates the isotropic two-spin exchange interactions $J_{ij}$ between magnetic sites $i$ and $j$. However, depending on the material, additional terms and interactions – such as single-ion anisotropy, Dzyaloshinskii-Moriya, and Kitaev interactions – can be included to account for relativistic effects like spin-orbit coupling 17. To enhance predictive accuracy and better reproduce experimental results, further extensions to the model, such as quantum renormalization factors 18, 19 or higher-order magnetic interactions like four-spin terms from cyclic exchange 20, 21 may become necessary.
@@
Developing methods to reliably and accurately determine the magnetic interaction parameters (MIPs) is still an active research field. The experimental determination of MIPs is generally achieved by fitting the measured spin-wave (magnon) dispersions. However, the resulting Hamiltonian is not uniquely defined because it depends on assumptions made about the model, such as the number of couplings and the type of interactions. In addition, even for the nearest-neighbor isotropic exchange coupling, the values obtained from this method can vary significantly. This variation arises not only from differences in experimental techniques but also from the fitting procedure itself, which depends on the number of $J_{ij}$ parameters considered and the inclusion of additional interactions in the Heisenberg Hamiltonian. Alternatively, MIPs can be computed directly from first principles using several established methods 17, 22, including the total-energy differences ($\Delta E$) approach 23, 24, 25, 26, 27, 28, 29, the infinitesimal-rotations method (IRM) based on the magnetic force theorem (MFT) 30, 31, 32, and the spin-spiral method that fits spin-spiral energy dispersions using the generalized Bloch theorem 33, 34, 35. Once the Heisenberg Hamiltonian is parameterized using one of these methods, magnon dispersions can be computed via linear spin-wave theory (LSWT) 36, 37. These methods rely on different approximations and are implemented with various technical nuances in different electronic-structure codes. Moreover, they are often applied on top of different descriptions of the material’s ground state, leading to significant discrepancies in the computed MIPs, as shown in numerous studies (see, e.g., Refs. 24, 25, 38). Therefore, a systematic comparison of these approaches is essential, particularly when applied to the ground state described using the same level of theory. In this work, we focus on the $\Delta E$ and IRM methods, which are described in greater detail in the following, and contrast these with a direct calculation of the spin-spin susceptibility as obtained from Hubbard-corrected time-dependent density-functional perturbation theory (TDDFPT+UU39.
@@
Two widely used flavors of the $\Delta E$ approach are prevalent in the literature. The first, and historically the earlier method, involves determining MIPs by computing the total energies of different spin configurations in supercells and then solving a system of linear equations to extract MIPs from the total energy differences 23, 24, 25. This approach requires calculating the total energy for $N+1$ spin configurations, which differ by the flipping of spins of some sites with respect to the ground state, with $N$ being the number of distinct MIPs. This method assumes that the exchange interactions are independent of spin configurations, which might not hold for all materials. Additionally, this approach provides quasi-averaged values of the microscopic interactions 29, and the self-consistency convergence can be challenging for some spin configurations. The second method is the four-state mapping analysis (FSMA) 26, 27, 28, 29. In this approach, a supercell is constructed and four spin states ($\uparrow\uparrow$, $\uparrow\downarrow$, $\downarrow\uparrow$, $\downarrow\downarrow$) are considered, each representing a different spin alignment of the spin pair whose interaction needs to be determined. A key advantage of the FSMA method is its ability to directly compute the derivatives of exchange interactions, thereby enabling the determination of spin-lattice coupling 27. However, this method is very sensitive to the supercell size used 40, 41, and thus the convergence of the MIPs with respect to the supercell size must be carefully verified.
@@
The IRM is based on MFT and leads to the Liechtenstein-Katsnelson-Antropov-Gubanov formula 30, 31, 32. It involves applying an infinitesimal spin rotation to the magnetic ground state and approximating the total energy change by the change in single-particle energies. IRM assumes that magnetic moments are localized on atoms, and it can only be applied when the system’s Hamiltonian is defined in a basis set of localized orbitals. This method can be implemented using various strategies, such as the Korringa–Kohn–Rostoker Green’s function method 42, 43, the tight-binding linear muffin-tin orbital method 44, 45, density-functional theory (DFT) 46, 47 with localized basis sets 48, 49, or DFT with plane-wave basis sets 50, 51 with subsequent wannierization to construct a tight-binding Hamiltonian using maximally localized Wannier functions (MLWFs) 52, 53. This latter technique is attractive due to the widespread use of plane-wave basis sets in electronic-structure codes. However, a drawback is the high sensitivity of the resulting interaction parameters to the quality of the atom-centering of the MLWFs. Depending on the system, MLWFs may not be perfectly centered on atoms, leading to significant variations in the computed MIPs 51, 49 and loss of symmetry. Furthermore, MIPs can be highly sensitive to various technical details in the IRM implementation 50, 49, 54, 55. Although IRM is widely used for computing MIPs 56, 57, 58, 24, its accuracy has been recently questioned 54, 55. Hence, assessing the accuracy of IRM relative to other methods is important.
@@
An alternative approach to determining MIPs is to compute magnon dispersion directly from the dynamical spin susceptibility and fit it using the analytic expression from LSWT. The dynamical spin susceptibility can be calculated using either time-dependent density-functional theory (TDDFT) 59 or many-body perturbation theory (MBPT) 60. TDDFT is typically used in the linear-response regime in the frequency domain, assuming a small external magnetic perturbation. Several techniques are available for this purpose, including the Dyson 61, 62, 63, 64, 65, 66, 67, Sternheimer 68, 69, 70, and Liouville-Lanczos 71 methods. To model ultrafast spin dynamics or strong perturbations, real-time propagation within TDDFT can also be employed 72. MBPT offers another route, based on solving the Bethe–Salpeter equation on top of a DFT or $GW$ ground state 73, 74, 75, 76, 77. While this approach is accurate, it is significantly more computationally demanding than TDDFT, especially when the latter uses the (computationally simple) adiabatic local spin-density approximation (ALSDA) kernel. However, ALSDA often fails to capture magnon excitations accurately in insulating transition-metal and rare-earth compounds. To address this limitation, more advanced TDDFT exchange-correlation kernels have been developed. Among them, ALSDA+$U$ 70, 78, 79, 39 has shown great promise. In this method, the Hubbard $U$ parameter – introduced within the DFT+$U$ framework 80, 81, 82, 83, 84 – plays a critical role, as the computed magnon energies are highly sensitive to its value. A particularly appealing strategy is the TDDFPT+$U$ approach of Ref. 39, which uses TDDFPT 71 in combination with ab initio Hubbard $U$ values. These $U$ parameters are computed self-consistently using linear-response theory 85, reformulated within the framework of density-functional perturbation theory (DFPT) 86, 87, 88. This method has demonstrated high accuracy across a wide range of materials 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, and holds strong potential for accurately and efficiently describing magnetic excitations from first principles.
@@
Various methods for directly calculating MIPs have been applied to transition-metal compounds such as NiO and MnO, with some examples found in Refs. 105, 24, 25, 38, 35, 106, 107, 108. However, comparing the computed $J_{ij}$ parameters across these studies is very challenging due to the use of different exchange-correlation functionals and corrective methods to describe the ground state. These include, in particular, self-interaction correction (SIC) methods 24, 25, hybrid functionals 25, and DFT+$U$ with either empirical or ab initio $U$ values 105, 35, 106, 38, 107, 108. The strong sensitivity of $J_{ij}$ to the value of $U$ further complicates comparisons between studies that used different $U$ values. Additionally, except for the study in Ref. 35, most research on NiO and MnO has focused on averaged isotropic exchange parameters $J_{ij}$, which neglects the splitting of nearest-neighbor $J_{ij}$ due to the rhombohedral distortion of the lattice. Moreover, no study has systematically compared MIPs from direct methods with those determined indirectly through LSWT fitting of magnon dispersion computed via dynamical spin susceptibility, such as using TDDFPT+$U$. Therefore, a comprehensive comparison of $J_{ij}$ values for transition-metal compounds, computed using both direct and indirect methods on top of an identical ground state and fully from first principles, is currently lacking.
@@
In this study, we perform a systematic comparison of the isotropic exchange interaction parameters $J_{ij}$ and magnon dispersions in NiO and MnO using three first-principles methods. The first two are direct methods: the $\Delta E$ method based on FSMA 26, 27, 28, 29, and the IRM based on MFT 30, 31, 32, implemented using a plane-wave basis and wannierization 50, 51. The $J_{ij}$ values obtained from both methods are then used to compute magnon dispersions via LSWT. The third approach is indirect: the TDDFPT+ $U$ method 39, based on the Liouville-Lanczos formalism 71. This method computes the dynamical spin susceptibility from first principles, from which the magnon dispersion is obtained directly and then fitted using LSWT to extract the $J_{ij}$ parameters. All three methods are applied consistently on top of the same DFT+$U$ ground state, with the Hubbard $U$ parameter computed ab initio using DFPT 86, 87, 88. The rhombohedral distortions of NiO and MnO are accounted for through DFT+$U$ structural optimizations. These distortions lead to a splitting of the nearest-neighbor exchange interactions for parallel and antiparallel spin pairs due to magnetoelastic coupling. Our results show that both the TDDFPT+$U$ method and the LSWT approach using $J_{ij}$ values from the $\Delta E$ method yield magnon dispersions in good agreement with inelastic neutron scattering data. In particular, both approaches accurately reproduce the finite magnon energy at the M point in the Brillouin zone (BZ), especially for MnO. In contrast, the LSWT results based on $J_{ij}$ values from the MFT method exhibit less accurate agreement with experiments, which may be attributed to implementation-specific details of this approach 49.
RESULTS
Magnetic order and Heisenberg Hamiltonians for NiO and MnO
Figure 1: Face-centered cubic conventional unit cell of NiO and MnO with the AFII magnetic order. Transition-metal elements (Ni in NiO, or Mn in MnO) and O atoms are shown in purple and grey, respectively. Yellow and blue vertical thick arrows centered on atoms indicate the direction of the spin. The ferromagnetic (111) planes are depicted in transparent blue and yellow color. The Heisenberg exchange interaction parameters $J_{1}^{+}$, $J_{1}^{-}$, and $J_{2}$ are highlighted for selected atoms, with thin red arrows showing the atom pairs involved in these interactions. Rendered using VESTA 109.

### v2

@@ I Introduction
Magnetic materials play a pivotal role in a wide range of modern technologies, from data storage and electronic devices to renewable energy solutions. They hold further promises in spintronics, magnonics, and quantum information [1, 2]. Noteworthy classes of magnetic materials that have drawn attention include high-temperature superconductors [3, 4], multiferroics [5, 6], 2D magnetic materials [7, 8], magnetic topological materials [9, 10], and, recently, altermagnets [11, 12, 13]. Various experimental techniques, such as inelastic neutron scattering, resonant inelastic X-ray scattering, and spin-polarized electron energy loss spectroscopy are employed to study these materials. Yet, theoretical models are essential for the interpretation of these experiments. Spin model Hamiltonians, such as the Ising, XY, and Heisenberg models [14] are powerful tools for understanding the microscopic origin of macroscopic properties and describing collective spin excitations and their coupling to phonons [15] and plasmons [16]. The Heisenberg model typically incorporates the isotropic two-spin exchange interactions $J_{ij}$ between magnetic sites $i$ and $j$. However, depending on the material, additional terms and interactions – such as single-ion anisotropy, Dzyaloshinskii-Moriya, and Kitaev interactions – can be included to account for relativistic effects like spin-orbit coupling [17]. To enhance predictive accuracy and better reproduce experimental results, further extensions to the model, such as quantum renormalization factors [18, 19] or higher-order magnetic interactions like four-spin terms from cyclic exchange [20, 21] may become necessary.
@@
Developing methods to reliably and accurately determine the magnetic interaction parameters (MIPs) is still an active research field. The experimental determination of MIPs is generally achieved by fitting the measured spin-wave (magnon) dispersions. However, the resulting Hamiltonian is not uniquely defined because it depends on assumptions made about the model, such as the number of couplings and the type of interactions. In addition, even for the nearest-neighbor isotropic exchange coupling, the values obtained from this method can vary significantly. This variation arises not only from differences in experimental techniques but also from the fitting procedure itself, which depends on the number of $J_{ij}$ parameters considered and the inclusion of additional interactions in the Heisenberg Hamiltonian. Alternatively, MIPs can be computed directly from first principles using several established methods [17, 22], including the total-energy differences ($\Delta E$) approach [23, 24, 25, 26, 27, 28, 29], the infinitesimal-rotations method (IRM) based on the magnetic force theorem (MFT) [30, 31, 32], and the spin-spiral method that fits spin-spiral energy dispersions using the generalized Bloch theorem [33, 34, 35]. Once the Heisenberg Hamiltonian is parametrized using one of these methods, magnon dispersions can be computed via linear spin-wave theory (LSWT) [36, 37]. These methods rely on different approximations and are implemented with various technical nuances in different electronic-structure codes. Moreover, they are often applied on top of different descriptions of the material’s ground state, leading to significant discrepancies in the computed MIPs, as shown in numerous studies (see, e.g., Refs. [24, 25, 38]). Therefore, a systematic comparison of these approaches is essential, particularly when applied to the ground state described using the same level of theory. In this work, we focus on the $\Delta E$ and IRM methods, which are described in greater detail in the following, and contrast these with a direct calculation of the spin-spin susceptibility as obtained from Hubbard-corrected time-dependent density-functional perturbation theory (TDDFPT+UU[39].
@@
Two widely used flavors of the $\Delta E$ approach are prevalent in the literature. The first, and historically the earlier method, involves determining MIPs by computing the total energies of different spin configurations in supercells and then solving a system of linear equations to extract MIPs from the total energy differences [23, 24, 25]. This approach requires calculating the total energy for $N+1$ spin configurations, which differ by the flipping of spins of some sites with respect to the ground state, with $N$ being the number of distinct MIPs. This method assumes that the exchange interactions are independent of spin configurations, which might not hold for all materials. Additionally, this approach provides quasi-averaged values of the microscopic interactions [29], and the self-consistency convergence can be challenging for some spin configurations. The second method is the four-state mapping analysis (FSMA) [26, 27, 28, 29]. In this approach, a supercell is constructed and four spin states ($\uparrow\uparrow$, $\uparrow\downarrow$, $\downarrow\uparrow$, $\downarrow\downarrow$) are considered, each representing a different spin alignment of the spin pair whose interaction needs to be determined. A key advantage of the FSMA method is its ability to directly compute the derivatives of exchange interactions, thereby enabling the determination of spin-lattice coupling [27]. However, this method is very sensitive to the supercell size used [40, 41], and thus the convergence of the MIPs with respect to the supercell size must be carefully verified.
@@
The IRM is based on MFT and leads to the Liechtenstein-Katsnelson-Antropov-Gubanov formula [30, 31, 32]. It involves applying an infinitesimal spin rotation to the magnetic ground state and approximating the total energy change by the change in single-particle energies. IRM assumes that magnetic moments are localized on atoms, and it can only be applied when the system’s Hamiltonian is defined in a basis set of localized orbitals. This method can be implemented using various strategies, such as the Korringa-Kohn-Rostoker (KKR) Green’s function method [42, 43], the tight-binding linear muffin-tin orbital (LMTO) method [44, 45], density-functional theory (DFT) [46, 47] with localized basis sets [48, 49], or DFT with plane-wave basis sets [50, 51] with subsequent wannierization to construct a tight-binding Hamiltonian using maximally localized Wannier functions (MLWFs) [52, 53]. This latter technique is attractive due to the widespread use of plane-wave basis sets in electronic-structure codes. However, a drawback is the high sensitivity of the resulting interaction parameters to the quality of the atom-centering of the MLWFs. Depending on the system, MLWFs may not be perfectly centered on atoms, leading to significant variations in the computed MIPs [51, 49] and loss of symmetry. Furthermore, MIPs can be highly sensitive to various technical details in the IRM implementation [50, 49, 54, 55]. Although IRM is widely used for computing MIPs [56, 57, 58, 24], its accuracy has been recently questioned [54, 55]. Hence, assessing the accuracy of IRM relative to other methods is important.
@@
An alternative approach to determining MIPs is to compute magnon dispersion directly from the dynamical spin susceptibility and fit it using the analytic expression from LSWT. The dynamical spin susceptibility can be calculated using either time-dependent density-functional theory (TDDFT) [59] or many-body perturbation theory (MBPT) [60]. TDDFT is typically used in the linear-response regime in the frequency domain, assuming a small external magnetic perturbation. Several techniques are available for this purpose, including the Dyson [61, 62, 63, 64, 65, 66, 67], Sternheimer [68, 69, 70], and Liouville-Lanczos [71] methods. To model ultrafast spin dynamics or strong perturbations, real-time propagation within TDDFT is employed [72]. MBPT offers another route, based on solving the Bethe–Salpeter equation on top of a DFT or $GW$ ground state [73, 74, 75, 76, 77]. While this approach is accurate, it is significantly more computationally demanding than TDDFT, especially when the latter uses the (computationally simple) adiabatic local spin-density approximation (ALSDA) kernel. However, ALSDA often fails to capture magnon excitations accurately in insulating transition-metal and rare-earth compounds. To address this limitation, more advanced TDDFT exchange-correlation kernels have been developed. Among them, ALSDA+$U$ [70, 78, 79, 39] has shown great promise. In this method, the Hubbard $U$ parameter – introduced within the DFT+$U$ framework [80, 81, 82, 83, 84] – plays a critical role, as the computed magnon energies are highly sensitive to its value. A particularly appealing strategy is the TDDFPT+$U$ approach of Ref. 39, which uses TDDFPT [71] in combination with ab initio Hubbard $U$ values. These $U$ parameters are computed self-consistently using linear-response theory [85], reformulated within the framework of density-functional perturbation theory (DFPT) [86, 87, 88]. This method has demonstrated high accuracy across a wide range of materials [89, 90, 91, 92, 93, 88, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103], and holds strong potential for accurately and efficiently describing magnetic excitations from first principles.
@@
The exchange interactions in NiO and MnO have been investigated from first-principles calculations in a number of prior works [104, 24, 25, 38, 35, 105, 106, 107]. These studies used different exchange-correlation functionals and corrective methods to describe the ground state, in particular, self-interaction correction (SIC) methods [24, 25], hybrid functionals [25], and DFT+$U$ with either empirical or ab initio $U$ values [104, 35, 105, 38, 106, 107]. These approaches differ in how they correct self-interaction errors, leading to variations in the resulting electronic structure, e.g. band gaps, magnetic moments, and charge densities. On top of these differing ground states, these studies employed various methods to compute $J_{ij}$ (e.g., MFT, $\Delta E$, or spin spiral), producing a spread of reported values with no clear consensus on either their absolute magnitudes or relative trends. Hence, comparisons of $J_{ij}$ are complicated due to the interplay of two factors: differences in the underlying ground-state description and differences in the computational approach used to determine the exchange parameters. Moreover, within the DFT+$U$ framework, the strong sensitivity of $J_{ij}$ to the value of $U$ further complicates comparisons between studies that used different $U$ values. Additionally, except for the study in Ref. 35, most research on NiO and MnO has focused on averaged isotropic exchange parameters $J_{ij}$, which neglects the splitting of nearest-neighbor $J_{ij}$ due to the rhombohedral distortion of the lattice. Moreover, no study has systematically compared MIPs from direct methods with those determined indirectly through LSWT fitting of magnon dispersion computed via dynamical spin susceptibility, such as using TDDFPT+$U$. Therefore, a comprehensive comparison of $J_{ij}$ values for transition-metal compounds, computed using both direct and indirect methods on top of an identical ground state and fully from first principles, is currently lacking.
@@
In this study, we perform a systematic comparison of the isotropic exchange interaction parameters $J_{ij}$ and magnon dispersions in NiO and MnO using three first-principles methods. The first two are direct methods: the $\Delta E$ method based on FSMA [26, 27, 28, 29], and the IRM based on MFT [30, 31, 32], implemented using a plane-wave basis and wannierization [50, 51]. The $J_{ij}$ values obtained from both methods are then used to compute magnon dispersions via LSWT. The third approach is indirect: the TDDFPT+ $U$ method [39], based on the Liouville-Lanczos formalism [71]. This method computes the dynamical spin susceptibility from first principles, from which the magnon dispersion is obtained directly and then fitted using LSWT to extract the $J_{ij}$ parameters. All three methods are applied consistently on top of the same DFT+$U$ ground state, with the Hubbard $U$ parameter computed ab initio using DFPT [86, 87, 88]. The rhombohedral distortions of NiO and MnO are accounted for through DFT+$U$ structural optimizations. These distortions lead to a splitting of the nearest-neighbor exchange interactions for parallel and antiparallel spin pairs due to magnetoelastic coupling. Our results show that both the TDDFPT+$U$ method and the LSWT approach using $J_{ij}$ values from the $\Delta E$ method yield magnon dispersions in good agreement with inelastic neutron scattering data. In particular, both approaches accurately reproduce the finite magnon energy at the M point in the Brillouin zone (BZ), especially for MnO. In contrast, the LSWT results based on $J_{ij}$ values from the MFT method exhibit less accurate agreement with experiments, which may be attributed to implementation-specific details of this approach [49].
II Computational details
All calculations are performed using Quantum ESPRESSO  (v7.2) [108, 109, 110]. Ground-state calculations are conducted with the PW code [108] using local spin-density approximation (LSDA) for the exchange-correlation functional including the Hubbard $U$ correction (LSDA+UU[111, 80]. We use norm-conserving scalar-relativistic pseudopotentials [112] from the PseudoDojo library [113]. For all calculations (unless otherwise stated), the plane-wave expansion of the Kohn-Sham wavefunctions is carried out with an 80 Ry kinetic-energy cutoff, and a 320 Ry cutoff is used for the charge density and potentials. The BZ is sampled using a $\Gamma$-centered $12\times 12\times 12$ $\mathbf{k}$ points grid for the 4-atom rhombohedral unit cell, unless stated otherwise, and type II antiferromagnetic (AFII) ordering is used (see Fig. 1). Spin-orbit coupling is neglected. Geometry optimization is performed using the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm [114], using a higher kinetic-energy cutoff of 90 Ry for the Kohn-Sham wavefunctions and 360 Ry for the charge density and potentials, with convergence criteria set to $10^{-8}$ Ry for total energy, $10^{-5}$ Ry/bohr for forces, and $0.01$ Kbar for pressure. All calculations are performed using the optimized LSDA+$U$ structural parameters. For NiO, the rhombohedral lattice parameter is $a=5.03$ Å and the rhombohedral angle is $\alpha=33.65^{\circ}$, while for MnO, $a=5.32$ Å and $\alpha=34.16^{\circ}$. The resulting magnetic moments and band gap values can be found in Ref. 39.
The Hubbard $U$ parameters are computed using DFPT [86, 88] (see Eq. (4) in the Supplemental Material (SM) [115]) as implemented in the HP code [116], with Löwdin-orthogonalized atomic orbitals for Hubbard projectors [117]. We employ uniform $\Gamma$-centered $\mathbf{k}$ and $\mathbf{q}$ points grids of size $8\times 8\times 8$ and $4\times 4\times 4$, respectively, for the 4-atom AFII unit cell, achieving an accuracy of approximately 0.01 eV for the Hubbard parameters. The $U$ parameters are computed iteratively in a self-consistent manner, as outlined in Ref. [87], and include Hubbard forces and stresses in LSDA+$U$ structural optimizations [118]. The final computed values for $U$ are 6.26 eV for the Ni($3d$) and 4.29 eV for the Mn($3d$) states in NiO and MnO, respectively.
The magnon energies are computed using TDDFPT+$U$ [39], as implemented in a turboMagnon code [119]. This method is based on computing the dynamical spin susceptibility using the Liouville-Lanczos approach [71] (see Eqs. (21) and (23) in the SM [115]). We use adiabatic LSDA+$U$ (ALSDA+$U$). The 4-atom AFII unit cell is used and the BZ is sampled with a $\Gamma$-centered $12\times 12\times 12$ $\mathbf{k}$ points grid. The calculations employ the pseudo-Hermitian flavor of the Lanczos recursive algorithm [120, 15], which includes an extrapolation technique for the Lanczos coefficients [121]. A total of 8000 Lanczos iterations are performed to achieve convergence in the TDDFPT+$U$ calculations. The TDDFPT+$U$ exchange parameters are derived by fitting the TDDFPT+$U$ magnon dispersions using LSWT (see Eqs. (7)–(9) in the SM [115]).
The exchange interaction parameters are computed using MFT (see Eqs. (18) in the SM [115]) as implemented in the TB2J code (v0.7.7.2) [51]. Calculations are performed with the 4-atom AFII unit cell. MLWFs are generated using the Wannier90 code (v3.1.0) [122], based on the LSDA+$U$ ground state. Specifically, we generate MLWFs for Mn($3d$), Mn($4s$), and O($2p$) states in MnO, as well as for Ni($3d$), Ni($4s$), and O($2p$) states in NiO, resulting in a total of 18 MLWFs out of 45 considered Kohn-Sham states for each material. We confirm that the Wannier centers for the Mn($3d$) and Ni($3d$) states are accurately centered on the atoms. For the ground-state calculations, the MLWFs generation, and the calculation of $J_{ij}$, the BZ is sampled using a $\Gamma$-centered $18\times 18\times 18$ $\mathbf{k}$ points grid. This sampling ensures that the calculation of the $J_{ij}$ parameters converges with an accuracy of 0.01 meV.
To compute the exchange interaction parameters using the $\Delta E$ method, we have developed a workflow within the AiiDA package [123], which will be described in detail elsewhere. We start with the AFII ground state in the LSDA+$U$ framework for the 4-atom unit cell, and then construct supercells of size $4\times 4\times 4$. The BZ is sampled using a $\Gamma$-centered $4\times 4\times 4$ $\mathbf{k}$ points grid for NiO and a $3\times 3\times 3$ $\mathbf{k}$ points grid for MnO. The plane-wave expansion of the Kohn-Sham wavefunctions and charge density employs kinetic-energy cutoffs of 80 Ry and 320 Ry for NiO, while for MnO we used 100 Ry and 400 Ry, respectively. This ensures the convergence of the $J_{ij}$ parameters with an accuracy of 0.01 meV for both materials. In these supercells, we initialize the original AFII configuration and then consider a target pair of spins in the following configurations: $\uparrow\uparrow$ (state 1), $\uparrow\downarrow$ (state 2), $\downarrow\uparrow$ (state 3), and $\downarrow\downarrow$ (state 4). For each configuration, we compute the total energy and then evaluate the exchange interaction parameters [see Eq. (13) in the SM [115]]. This procedure allows us to determine the exchange parameters $J_{ij}$ depending on the selected target pair of spins in the supercell.
The magnon dispersions based on LSWT, using the exchange interaction parameters reported in Table 1, are computed with the SpinW code (v3.1) [124]. The fitting of all magnon dispersions in Fig. 2 is performed using the minimal model [see Eq. (2)] and the analytic LSWT expressions [see Eqs. (7)–(9) in the SM [115]].
The data used to produce the results of this work is available in the Materials Cloud Archive [125].
III RESULTS
III.1 Magnetic order and Heisenberg Hamiltonians for NiO and MnO